OpenStudy (anonymous):

Simplify the following expression. 3x − 4y + 6z − x − 3y + 2z I got 3x-1y+8z. I subtracted 4y and 3y, thats where i got my 1y. Then i added 6z and 2z, which i got 8z. I looked at the answer choices and i couldn't find that as a choice. Here are the choices: a: 4x − y + 8z b: 2x − y + 8z c: 2x − 7y + 8z d: 2x − 7y + 12z

OpenStudy (zale101):

Collect the like terms for the first question

OpenStudy (zale101):

3x − 4y + 6z − x − 3y + 2z 3x-x = -4y-3y= 6z+2Z=

OpenStudy (zale101):

what i did in here is that i combined the like terms. X with x, y with y, and z with z

OpenStudy (zale101):

3x-x =? -4y-3y=? 6z+2z=?

OpenStudy (anonymous):

I never noticed the x. I saw the 3x but not the single x. I think that threw me off. @Zale101

OpenStudy (zale101):

the single x basically stands for 1

OpenStudy (anonymous):

4x-y+8z ? @Zale101

OpenStudy (zale101):

no

OpenStudy (zale101):

how did u solve that?

OpenStudy (zale101):

3x-x is 2x -4y-3y is -7y 6z+2z is 8z

OpenStudy (zale101):

2x-7y+8z is the right answer, not 4x-y+8z :)

OpenStudy (zale101):

@Puurfect_kitty , u there?

OpenStudy (anonymous):

yes, sorry, i skipped that question to the ones i knew. and i did it like i added 3x and x. (4x). I subtracted 4y and -3y (-1y). then added 6z and 2z (8z) @Zale101

OpenStudy (zale101):

the mistake u did is that 4y supposed to be -4y 3x-x, subtract 3x - 4y + 6z - x -3y + 2z 2x - 7y + 8z

OpenStudy (anonymous):

Ok, thank you for the help(: @Zale101

OpenStudy (zale101):

i'm glad it did :D